OpenStudy (anonymous):

Identify the variation as direct, inverse, joint or combined. y = 7x direct inverse joint combined

OpenStudy (anonymous):

@MasterprojectHD

OpenStudy (anonymous):

inverse

OpenStudy (anonymous):

wait no yeah its inverse

OpenStudy (anonymous):

Yep its inverse

OpenStudy (anonymous):

Identify the variation as direct, inverse, joint or combined. xy = 24

OpenStudy (anonymous):

what do u think it is

OpenStudy (anonymous):

inverse

OpenStudy (anonymous):

Inverse variation.

OpenStudy (anonymous):

Identify the variation as direct, inverse, joint or combined. F=4/R

OpenStudy (anonymous):

Direct variation.

OpenStudy (anonymous):

Identify the variation as direct, inverse, joint or combined. m/n=4

OpenStudy (anonymous):

Direct

OpenStudy (anonymous):

Identify the variation as direct, inverse, joint or combined. i = 2pr

OpenStudy (anonymous):

Joint

OpenStudy (anonymous):

Identify the variation as direct, inverse, joint or combined. \[R=cl/r^2\]

OpenStudy (anonymous):

Combined

OpenStudy (anonymous):

Identify the variation as direct, inverse, joint or combined. The cost of the number of the same articles purchased varying with the number purchased.

OpenStudy (anonymous):

Direct.

OpenStudy (anonymous):

No I still need a lot of help.

OpenStudy (anonymous):

i gtg :(

OpenStudy (anonymous):

Noooooo

OpenStudy (anonymous):

How many do you have left?

OpenStudy (anonymous):

12 or 13

OpenStudy (anonymous):

ok ill try

OpenStudy (anonymous):

Identify the variation as direct, inverse, joint or combined. If the area of a rectangle stays constant, what is the relationship between the base and the height?

OpenStudy (anonymous):

Combined

OpenStudy (anonymous):

Identify the variation as direct, inverse, joint or combined. A = pi(r²)

OpenStudy (anonymous):

Inverse

OpenStudy (anonymous):

Identify the variation as direct, inverse, joint or combined. V = pi(r²h)

OpenStudy (anonymous):

Im not sure about the last one. A = pi(r^2) but I think its inverse

OpenStudy (anonymous):

Which one?

OpenStudy (anonymous):

@MasterprojectHD